### v1

# Terrain Diffusion: A Diffusion-Based Successor to Perlin Noise in Infinite, Real-Time Terrain Generation

Abstract.
For decades, procedural worlds have been built on procedural noise functions such as Perlin noise, which are fast and infinite, yet fundamentally limited in realism and large-scale coherence. We introduce Terrain Diffusion, an AI-era successor to Perlin noise that bridges the fidelity of diffusion models with the properties that made procedural noise indispensable: seamless infinite extent, seed-consistency, and constant-time random access. At its core is InfiniteDiffusion, a novel algorithm for infinite generation, enabling seamless, real-time synthesis of boundless landscapes. A hierarchical stack of diffusion models couples planetary context with local detail, while a compact Laplacian encoding stabilizes outputs across Earth-scale dynamic ranges. An open-source infinite-tensor framework supports constant-memory manipulation of unbounded tensors, and few-step consistency distillation enables efficient generation. Together, these components establish diffusion models as a practical foundation for procedural world generation, capable of synthesizing entire planets coherently, controllably, and without limits.
copyright: none
Figure 1. A region of a world generated with Terrain Diffusion. The leftmost panel spans roughly five million square kilometers, with about 2.2 million square kilometers of land area, comparable to the size of the Congo. Red boxes denote the region shown in the next panel, illustrating coherent terrain generation across four orders of magnitude in scale. Zoom for details.: :
1. Introduction
Procedural terrain generation underpins the creation of virtual worlds, from open-world games to planetary simulations. For nearly four decades, procedural noise functions such as Perlin noise (Perlin, 1985, 2002) have defined this field. They offer three properties that make them indispensable for procedural worlds: seamless infinite extensibility, seed-consistency, and constant-time random access. A single random seed can deterministically produce a boundless landscape without storing vast datasets, providing an elegant foundation for procedural worlds.
Yet these procedural methods are inherently limited. Their patterns are smooth and lack the hierarchical organization of real geography. Continents, mountain ranges, and river basins emerge in nature from structured, multi-scale processes that simple noise cannot capture. As a result, worlds built from procedural noise often appear plausible but not real.
Recent advances in generative modeling, particularly diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), have transformed image synthesis by learning to reproduce natural structure with remarkable realism and control, but these methods are typically confined to smaller bounded domains. Recent work has explored infinite or large-scale generation capabilities, but these approaches generally lose one or more of the core properties that make procedural noise valuable in interactive applications.
Terrain Diffusion bridges this divide, introducing a learned successor to procedural noise that preserves all its core properties: seamless infinite extensibility, seed-consistency, and constant-time random access. To do so, we generalize MultiDiffusion (Bar-Tal et al., 2023) for infinite inference, enabling real-time, coherent, procedural generation.
A hierarchical diffusion stack unifies global and local structure: a coarse planetary model establishes continental structure, refined by a higher-resolution model that introduces mountain ranges, valleys, and local relief. A novel elevation encoding further stabilizes training and inference across the full dynamic range of Earth’s terrain. Few-step consistency distillation (Song et al., 2023; Lu and Song, 2025) enables real-time inference, and an open-source infinite-tensor framework enables constant-memory streaming and composable manipulation of infinite tensors.
Together, these components establish the first learned system capable of streaming an entire planet in real time on consumer GPUs. Terrain Diffusion demonstrates that diffusion models can serve as a practical foundation for infinite, seed-consistent worlds that can be explored interactively, in real-time, and without restrictions.
2. Related Works
Procedural noise. Procedural terrain traditionally relies on procedural noise such as Perlin or Simplex noise (Perlin, 1985, 2002), often combined with fractal Brownian motion (fBm) (Fournier et al., 1982). These methods remain popular for their controllability, speed, seed-consistency, and infinite tileability, but they lack the large-scale structure of real landscapes. They produce repetitive, texture-like patterns and cannot reproduce complex features such as mountain ranges, branching valleys, or volcanic and glacial landforms without extensive post-processing.
Diffusion and consistency models. Denoising diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015) generate samples by iteratively refining noise and are widely used for high-fidelity synthesis. Consistency models (Song et al., 2023) approximate denoised diffusion outputs in one or a few steps, and continuous variants (Lu and Song, 2025) achieve throughput competitive with GANs while retaining most of the quality of full diffusion sampling, enabling real-time use cases.
Learned terrain generation. GAN-based terrain models (Goodfellow et al., 2020; Voulgaris et al., 2021; Spick and Walker, 2019; Beckham and Pal, 2017) can generate convincing local relief, but they operate on fixed crops and do not tile, limiting them to bounded worlds. Diffusion-based terrain and Earth imagery synthesis (Hu et al., 2024; Guérin et al., 2017; Borne-Pons et al., 2025) further improves fidelity and control, but also assumes finite canvases and requires relatively significant compute. Sharma et al. (Sharma et al., 2024) achieves $1024{\times}$ super-resolution with strong global coherence, yet still targets bounded images, does not run in real time, and does not generate elevation. Jain et al. (Jain et al., 2023) is closest to our work, offering infinite terrain generation by sampling diffusion-based tiles and blending them with a Perlin-based kernel. Because tiles are generated independently and the kernel has no awareness of broader context, global coherence and large scale structure remain tied to procedural noise rather than the learned model. In contrast, Terrain Diffusion couples all tiles through a shared global context and fuses tiles through a fully learned, context-aware mechanism. Procedural noise is used only for continental-scale structure, where data is sparse and simple enough that more complex alternatives would provide little benefit while reducing user control.
MultiDiffusion and unbounded generation. Several works extend generative models beyond image bounds seen in training. InfinityGAN (Lin et al., 2022) produces infinite images with GANs but does not extend to diffusion models, limiting scalability. MultiDiffusion (Bar-Tal et al., 2023) and Mixture of Diffusers (Jiménez, 2023) generate images larger than the training canvas but still assume a bounded final extent. BlockFusion (Wu et al., 2024) and WorldGrow (Li et al., 2025) generate worlds by conditioning each tile on its neighbors, producing continuous worlds but without seed consistency, since outputs depend on sampling order. In contrast, Terrain Diffusion defines a seed-consistent InfiniteDiffusion algorithm whose outputs are order invariant and allow constant-time random-access generation over an infinite domain.
3. InfiniteDiffusion: Unbounded Generation Across Planetary Scales
MultiDiffusion (Bar-Tal et al., 2023) provides a simple and effective way to extend diffusion sampling beyond a model’s native resolution by averaging predictions from overlapping windows. This enables synthesis across larger images, as local predictions fuse into a seamless image. However, in its standard form, MultiDiffusion remains confined to bounded domains: all windows must lie within a fixed finite canvas, limiting its applicability to unbounded worlds or continuously streamed environments.
We introduce InfiniteDiffusion, an extension of MultiDiffusion that lifts this constraint. By reformulating the sampling process to operate over an effectively infinite domain, InfiniteDiffusion supports seamless, consistent generation at scale. The remainder of this section reviews the principles of MultiDiffusion, and formalizes its extension to unbounded domains. We present the definitions in $\mathbb{Z}^{2}$ for clarity, but all results extend to $\mathbb{Z}^{d}$ with minimal modification.
3.1. A Review of MultiDiffusion
MultiDiffusion extends standard diffusion sampling by averaging overlapping windows, producing consistent outputs from local predictions. Each denoising step aggregates the predictions of overlapping patches, enforcing continuity across window boundaries and allowing generation of regions much larger than a model’s input size. The process begins with a pretrained diffusion model $\Phi$, operating on images in $\mathcal{I}=\mathbb{R}^{H\times W\times C}$. The diffusion process generates a sequence of images
$I_{T},I_{T-1},\dots,I_{0}\quad\text{s.t.}\quad I_{t-1}=\Phi(I_{t}\mid y)$
that refines the original noisy image $I_{T}$ into a fully denoised version $I_{0}$, under conditioning vector $y$. MultiDiffusion defines a new model $\Psi$ that generates in a different image space $\mathcal{J}=\mathbb{R}^{H^{\prime}\times W^{\prime}\times C}$, producing a new sequence of images
$J_{T},J_{T-1},\dots,J_{0}\quad\text{s.t.}\quad J_{t-1}=\Psi(J_{t}\mid z).$
To accomplish this, MultiDiffusion defines $n$ windows indexed by $i\in[n]$. In the finite setting, a region $R$ is any rectangular subset of the $H^{\prime}\times W^{\prime}$ coordinate grid, while a window region has a fixed size $H\times W$. In MultiDiffusion, each window $i$ is assigned a window region $R_{i}$. For an image $J\in\mathcal{J}$, we write $J[R]\in\mathcal{I}$ for the values of $J$ on the coordinates in $R$.
Each window also has a weight matrix $W_{i}\in\mathbb{R}^{H\times W}$ that specifies the relative contribution of each pixel in the pretrained diffusion model’s output. Let $U_{i}(x)$ denote the $H^{\prime}\times W^{\prime}$ image that places an $H\times W$ tensor $x$ in the region $R_{i}$ with zeros elsewhere. With these definitions, the closed form MultiDiffusion update for direct pixel or latent-space samples is
(1) $\Psi(J_{t}\mid z)=\frac{\sum_{i=1}^{n}U_{i}(W_{i}\otimes\Phi(J_{t}[R_{i}]\mid y_{i}))}{\sum_{j=1}^{n}U_{j}(W_{j})}$
where $\otimes$ denotes the Hadamard product. This expression represents a weighted average of all local denoising predictions, where each window contributes according to its weight map. The result is a global update that reconciles all overlapping diffusion paths into a single image.
Although MultiDiffusion elegantly unifies local diffusion paths, it remains constrained to bounded domains: the process assumes a finite number of windows and requires the pretrained diffusion model to be evaluated at all windows to complete one step. Extending the same principle to infinite domains therefore requires reformulating $\Psi$ so that it operates locally and independently of global window layouts, a key step towards the InfiniteDiffusion algorithm introduced next.
3.2. From MultiDiffusion to InfiniteDiffusion
We now seek to extend MultiDiffusion beyond finite image domains. We first redefine the MultiDiffusion image space as an unbounded image, $\mathcal{J}=\mathbb{R}^{\mathbb{Z}\times\mathbb{Z}\times C}$, so that generation produces an infinite output. Consequently, we now define a region to be any rectangular subset of $\mathbb{Z}^{2}$. Since generation is now over an infinite domain, window indices must now range over a countably infinite set $S$.
For all applications shown in this work, we take $S=\mathbb{Z}^{2}$, so each window is indexed by $(i,j)$, and each window region $R_{ij}$ is defined as a square sliding window with side length $H=W$ and stride $s$ on both axes. Concretely, $R_{ij}=[is,\,is+H)\times[js,\,js+W)$. This particular layout is not essential to the InfiniteDiffusion formulation and serves only as an implementation choice for the experiments.
With an infinite number of windows, the MultiDiffusion update becomes intractable, and computation requires an infinite sum to produce the final image. Instead, we seek to generate the image lazily, by only evaluating particular regions $R$.
To achieve this, we define $\kappa$ to be the function mapping a region to the set of window indices that overlap it. We assume $|\kappa(R)|$ is always finite. This enables the InfiniteDiffusion update
(2) $\Psi(J_{t}\mid z)[R]=\left(\frac{\sum_{i\in\kappa(R)}U_{i}(W_{i}\otimes\Phi(J_{t}[R_{i}]\mid y_{i}))}{\sum_{j\in\kappa(R)}U_{j}(W_{j})}\right)[R],$
which is the MultiDiffusion update with only the windows intersecting $R$ evaluated. In the finite setting, the full image $J_{t}$ can be generated in advance, making each $J_{t}[R_{i}]$ effectively free. In the infinite setting, precomputing $J_{t}$ is impossible, so evaluating $J_{t}[R_{i}]$ requires recursively invoking the same update. A direct implementation would therefore incur exponentially growing compute, motivating an alternative strategy.
Figure 2. A conceptual visualization of InfiniteDiffusion with sliding windows. The user’s query $J_{0}[R]$ induces a deterministic chain of window queries: computing $J_{0}[R]$ requires a region $J_{1}[R_{0}]$, which in turn requires a region $J_{2}[R_{1}]$. Querying $J_{2}[R_{1}]$ is inexpensive since it corresponds directly to Gaussian noise.:
3.3. Practical Querying of InfiniteDiffusion
To make queries practical, we avoid recomputing the same window updates across recursive calls. Instead, for each image $J_{t}$ we maintain two corresponding infinite tensors: $A_{t}$, which stores the numerator in Eq.  2, and $B_{t}$, which stores the denominator. Both tensors are initialized to zero. These infinite tensors are typically stored in tiles, which can be loaded and unloaded as necessary to conserve memory. When a query $J_{t}[R]$ occurs, we identify all the windows required to generate the region, and process all previously unprocessed windows, populating the desired regions of $A_{t}$ and $B_{t}$. Then $J_{t}[R]=A_{t}[R]/B_{t}[R]$. Final generation proceeds as a recursive process that begins by sampling $J_{T}$ as Gaussian noise. $J_{0}[R]$ is obtained by recursively applying the query routine at all earlier steps. In summary, the routine below computes $J_{t-1}[R]$ by evaluating only the windows that overlap $R$, caching each window’s contribution in $A_{t-1}$ and $B_{t-1}$ so that future queries can reuse the same results. The tensors $A_{t-1}$, $B_{t-1}$, and the set of processed windows $P$ are mutated in-place.
Algorithm 1 Querying $J_{t-1}[R]$ with InfiniteDiffusion. See Fig.  2
Inputs:
[…]
for each window $i$ in $\kappa(R)\setminus P$ do
  $A_{t-1}[R_{i}]\leftarrow A_{t-1}[R_{i}]+W_{i}\otimes\Phi(J_{t}[R_{i}]\mid y_{i})$
  $B_{t-1}[R_{i}]\leftarrow B_{t-1}[R_{i}]+W_{i}$
end for
$P\leftarrow P\cup\kappa(R)$
Output: $J_{t-1}[R]=A_{t-1}[R]/B_{t-1}[R]$
There is only one additional challenge in making InfiniteDiffusion practical in real-world applications. Each query $J_{t-1}[R]$ typically requires a region of $J_{t}$ larger than $R$ itself, as pixels near the edge of $R$ are generated by windows that extend beyond $R$. In the sliding-window case, regions grow quadratically in area, yielding an overall $O(T^{3})$ time complexity for final generation. While this prevents the use of longer noise schedules, we find that the method remains both fast and effective even when $T$ is near 1, provided each $y_{i}$ captures sufficient context. In practice, this is achieved by either using a few-step model, such as a consistency model, or simulating one by running several diffusion steps at a time.
3.4. Properties of InfiniteDiffusion
Formal proofs for all properties are provided in appendix   A
Seed consistency.
A central property of InfiniteDiffusion is that it preserves the seed-consistent behavior that makes procedural noise functions useful for procedural world generation. Seed consistency follows directly from how the process is defined. Once a seed is fixed, the initial noise image $J_{T}$ is completely determined. The query routine is a deterministic function of its inputs, so $J_{t-1}$ is a deterministic function of $J_{t}$, and caching only memoizes intermediate results of this deterministic computation. This argument requires that the conditioning variables $y_{i}$ are themselves seed-consistent, which is typically trivial when $y_{i}$ is constant or generated by the same InfiniteDiffusion algorithm, as in this work. By composing these steps, every image $J_{t}$ and every region $J_{t}[R]$ becomes a fixed function of the original seed. In particular, requesting the same region again, or in a different order relative to other regions, always returns exactly the same value, since the computation depends only on the seed and the region, not on the sequence of queries.
Constant-time random access.
Assuming that $|\kappa(R_{i})|\leq M$ for any window region $R_{i}$, InfiniteDiffusion guarantees constant-time random access. In particular, for any window region $R_{i}$, the query $J_{0}[R_{i}]$ is $O(1)$. In practice, this allows discontinuous exploration of the world: any region can be queried independently, and seed consistency ensures that skipping intermediate tiles never alters the generated content or its quality.
Parallelization.
InfiniteDiffusion also admits parallel evaluation of window updates. For any fixed timestep $t$, each evaluation of $\Phi$ is independent, so they can be batched and executed in parallel.
3.5. An Open Source Infinite Tensor Framework for Unbounded Inference
To support unbounded generation without exceeding memory limits, we introduce the Infinite Tensor framework, a Python library that enables sliding window computation over tensors with infinite dimensions. It allows models to process arbitrarily large images as if they were standard PyTorch tensors while keeping only the visible region in memory. Each operation is performed through a fixed-sized sliding window that dynamically loads and evicts data as sampling progresses, permitting inference on arbitrarily large scenes with constant memory use.
This abstraction lets diffusion and consistency models operate directly on infinite images without manual data management. Windows can overlap to provide context and blend results, and multiple infinite tensors can depend on one another to form hierarchical pipelines. The framework serves as the runtime layer that links local model inference with practical, global world synthesis.
Together, InfiniteDiffusion and the Infinite Tensor framework provide the foundation required for practical, unbounded generation. The remaining components of Terrain Diffusion build on this foundation by combining these capabilities with large-scale real-world training data, hierarchical modeling, and a task-specific architecture.
4. Data
Our dataset combines multiple global sources to provide consistent coverage of both land and ocean. Land elevations are drawn from the 3-arc-second MERIT DEM (Yamazaki et al., 2017), while ocean bathymetry is taken from the 30-arc-second ETOPO dataset (NOAA National Geophysical Data Center, 2009). Since ETOPO’s resolution is lower, it is blurred and upsampled to match MERIT’s resolution before merging. To ensure smooth coastal transitions, we measure the distance of each ocean pixel from the nearest coastline and linearly interpolate elevation from 0 m at the shore to the local ocean depth 100 pixels offshore. To support climatic conditioning, we supplement elevation with 30-arc-second WorldClim (Fick and Hijmans, 2017) data for temperature and precipitation.
For efficient processing, data is downloaded and stored in contiguous $2048\times 2048$ tiles, each covering equal surface area at approximately 90m resolution. To maintain uniform ground resolution, tiles are stretched in longitude so that each pixel represents roughly the same area regardless of latitude. This equal-area tiling ensures consistent scale across training and allows all models to train on data without distortion. Finally, 80% of the tiles are randomly assigned to the train set, and the remainder are withheld for validation.
All models are trained on random crops drawn from random tiles, with sampling biased so that 99% of tiles contain at least 1% land, as ocean regions are simpler and lower priority. Each crop is randomly flipped and rotated in 90 degree increments to reflect our goal of generating infinite, directionless terrain, similar in spirit to Perlin noise.
5. Hierarchical Modeling & Stabilization
This section outlines the hierarchical architecture and data representation underlying our pipeline, which together enable coherent, high-fidelity terrain generation across planetary scales.
Figure 3. Effects of the signed-sqrt transform. Standard deviation become more uniformly distributed with respect to mean elevation, and the range of standard deviations compress.
5.1. Signed Square-Root Transform
Terrain tiles vary significantly in elevation range. Under a fixed diffusion noise schedule, this leads to uneven effective SNR: low-relief regions behave as though exposed to stronger noise, while high-relief terrain is affected much less. To reduce this variation, we apply a signed square-root transform $z\mapsto\mathrm{sign}(z)\sqrt{|z|}$. Figure 3 illustrates its effect by comparing tile means and standard deviations before and after the transform. In the raw elevation space, tiles with higher absolute elevations exhibit larger variance. The transform compresses high-relief values and distributes variance across tiles more uniformly. In practice, this allows us to train with a more focused noise distribution and enhances the visibility of small features, especially coastlines.
(a) Initial Input
(b) Refined Coarse Map
(c) Final 90m Elevation
Figure 4. Multi-stage elevation generation pipeline. (a) The initial coarse map, which serves as the structural and climatic guide. Can be made by hand or generated procedurally. (b) The refined coarse map, enhanced by our lightweight coarse model to enforce realism. (c) The final 90m elevation map generated by the core latent diffusion model, with InfiniteDiffusion for tiling.
5.2. Stabilization Via Laplacian Encodings
Due to normalization, the large dynamic range of Earth elevations make model errors deceptively large in absolute units. Without the signed-sqrt transform, even relatively small errors of $\sigma=0.01$ can correspond to $\pm 25$m noise after denormalization. Applying the transform lessens the issue at lower elevations, but worsens it at higher ones. While this noise is less obvious in images, it can become very apparent in interactive settings, where the noise is potentially much larger than the viewer or the surrounding scenery. To mitigate this, we predict a Laplacian-based representation comprising a low-frequency component, obtained by downsampling and blurring the original image, and a residual/high-frequency component given by subtracting the upsampled low frequency component from the original image.
Residual errors are over $30\times$ smaller in magnitude due to their lower variance. To clean the low-frequency channel after generation, we decode the noisy low- and high-frequency components $(L+H)$ into a provisional heightmap, then blur and downsample it to re-extract a denoised low-frequency $\hat{L}$, with any high-frequency noise redirected to the residual $\hat{H}$. Final synthesis uses $\hat{L}+H$, so low-frequency errors vanish while high-frequency detail is preserved. In practice, $L\approx\hat{L}$ even under strong synthetic noise, confirming that re-extraction cleanly isolates low-frequency structure. This Laplacian denoising step reduces the FID of the untiled core diffusion model (introduced next) from $23.21$ to $9.33$, and the corresponding consistency model from $75.75$ to $12.71$.
5.3. A Hierarchical Model
Planetary terrain spans several orders of magnitude in scale, from continental structure to meter-level detail, making one-pass generation infeasible. Several previous works (Sharma et al., 2024; Zhang et al., 2025; Lee et al., 2023; Du et al., 2024; Zhou and Tang, 2024) have shown that MultiDiffusion produces incoherent and repetitive results when poorly conditioned, and InfiniteDiffusion does not natively provide a solution to this. We therefore organize generation into a small hierarchy of models operating at progressively finer resolutions. Each stage refines and conditions on the one above, maintaining planetary coherence while producing realistic local detail. All models share a common EDM2 (Karras et al., 2024b) backbone with the modifications proposed in sCM (Lu and Song, 2025). The hierarchy begins with a coarse planetary model, which generates the basic structure of the world from a rough, procedural or user-provided layout. The next stage is the core latent diffusion model, which transforms that structure into realistic 46km tiles in latent space. Finally, a consistency decoder expands these latents into a high-fidelity elevation map. We visualize the coarse-to-fine pipeline in Figure  4.
The core latent diffusion model synthesizes 512×512 patches at a 90m resolution in signed-sqrt space, corresponding to 46km tiles. It predicts a 64x64 low-frequency elevation channel and latent map that compactly represents the corresponding residual. To supply the latent codes, we train a separate VAE-style autoencoder that shares the same U-Net backbone but omits diffusion-specific components. The model is optimized using L1 and LPIPS losses with a weak KL term to prevent overfitting. After training, the encoder processes each 2048×2048 tile in the dataset as a whole, and the resulting latent image is precomputed and stored alongside the tile. To maximize local quality, the imprecise VAE decoder is discarded, and a diffusion decoder learns to expand these latents into realistic and high-resolution residuals. During training of all models, we draw random crops from the latent image to learn a nearly translation-invariant representation, reflecting the fact that generation should be independent of absolute location. Conditioning is implemented by concatenating nearest-neighbor interpolated latents to the noisy input image at each diffusion step. We train both the autoencoder and diffusion decoder on 128×128 crops, but find the models generalize well when applied to higher-resolution patches.
To facilitate long-range global coherence, the core model is conditioned on 4x4 patches of elevation data. Each pixel of the patch is about 23km, with the model prediction corresponding to the 2x2 interior. Each patch contains 3 channels: the mean elevation of the pixel, the 5th percentile elevation of the pixel, and a binary mask indicating which pixels have data available. We also provide the model with the tile’s mean temperature, temperature variation, annual precipitation, and precipitation seasonality for additional coherence and user control. Since climatic data is not available in the ocean, we replace missing values with a standard gaussian, ensuring the model accepts any combination of climatic variables in ocean regions.
Additionally, we optionally condition on a target distribution over tile ”beauty” scores. A linear regressor trained on 150 hand-labeled examples and hand-crafted Fourier-based features ($R^{2}\approx 0.8$) assigns each tile a beauty value; during training, we generate random score histograms, sample a random score from it, select an appropriate tile, and feed the target distribution as conditioning, which later allows users to bias synthesis toward more or less visually striking terrain. All conditioning data is provided as a flat tensor, and injected into the model alongside the noise embedding.
5.4. Real-Time Planetary Scale Synthesis
While the 46 km tiles are realistic in isolation, large-scale coherence requires conditioning on broader planetary context. To provide this, we introduce a compact coarse diffusion model that generates the channels required for conditioning the core diffusion model. The user provides initial maps for these variables using procedural noise or hand-drawn sketches. During inference, these inputs are corrupted with gaussian noise according to the user’s preference on a per-channel basis, and concatenated against the usual diffusion inputs. The model follows the EDM2 design but with no downsampling or upsampling operations. This limits the receptive field by design, preventing the model from drifting toward the massive continental structures present in Earth data and avoiding conflicts where global priors override user guidance while still allowing strong local corrections.
To support real-time synthesis, all diffusion models, except the coarse model, are distilled into continuous-time consistency models (Lu and Song, 2025). To improve fidelity further, we apply the guidance scheme proposed in AutoGuidance (Karras et al., 2024a). Combined, these stages form a continuous generation pipeline, from planetary context to local detail, capable of on-demand, real-time synthesis.
6. Results
We evaluate Terrain Diffusion on visual fidelity and latency. All experiments use a single NVIDIA RTX 3090 Ti GPU. InfiniteDiffusion uses a stride of 32 in the core model and a separable linear weight window that is 1 at the window center and decreases linearly to a small value $\epsilon$ at the boundary, which reduces FID from $25.83$ with a constant map to $17.87$. The decoder model uses windows of size 512 with strides 384. Batch size is 32 for the core model and 1 for other components.
6.1. Visual Fidelity
For quality, we report FID-50k for (1) non-tiled diffusion samples, (2) non-tiled consistency samples, and (3) tiled samples generated with InfiniteDiffusion [^1] FID is computed on central 984×984 crops of the validation tiles to ensure adequate surrounding context. InfiniteDiffusion is applied only to the latent space; tiling the decoder would require impractically large context, and the decoder already produces near seamless outputs.. This isolates base fidelity, the effect of consistency distillation, and the effect of tiling. Evaluations use raw elevation values, with images normalized by centering each tile and scaling by the larger of its value range or 255, ensuring that images are not expanded beyond the native precision of the data. All results use two-step generation for both the core consistency model and the consistency decoder. Table 1 summarizes our results.
For context, we also measure the decoder’s standalone rFID at $512\times 512$ resolution. The one step variant obtains an rFID of 3.08, while the two step variant reaches 1.07. The two step model uses a small intermediate noise level of $\sigma=0.065$, indicating that the second step performs only minor refinements that may be unnecessary for speed sensitive applications.
Table 1. FID-50k comparison for central crops generated with InfiniteDiffusion vs. direct generations. Lower is better.
Tiling Model Type FID $\downarrow$
None Diffusion 9.34
None Consistency 12.71
InfiniteDiffusion Consistency 17.87
6.2. Real Latency: Time to First and Second Tile
Because generation of any fixed-size region has bounded cost, generating a contiguous $512\times 512n$ strip is $O(n)$. But neighboring tiles reuse cached context, so the cost for querying the first $512\times 512$ region is larger than for subsequent regions.
Motivated by these facts, we measure latency as the time to first tile (TTFT) and time to second tile (TTST) across resolutions. TTFT denotes the delay from model initialization to the first $\times512$ tile becoming available, reflecting initial setup cost. TTST measures the time to generate an adjacent $512\times 512$ tile thereafter, reflecting interactive exploration performance. In practice, while both metrics are bounded, they vary with the specific region location because the number of intersecting windows differs across positions. To account for this, as well as minor variations between runs, we perform each evaluation 100 times at random locations and report the average and standard deviation.
Table 2. Generation latency for the first and second tile.
Metric Seconds
TTFT 7.60 ± 1.00
TTST 2.40 ± 1.06
An F-35, one of the fastest conventional aircraft in service at roughly 550 m/s, would traverse a 512×512 tile at 90 m resolution in about 84 seconds. In that time, Terrain Diffusion can produce 35 additional tiles, which places the latency well within the requirements of real-time streaming for any slower application. Even if the terrain is downsampled substantially, reducing traversal time by more than an order of magnitude, the query rate still remains well within the model’s capacity.
6.3. Qualitative Analysis
Figure 5 shows 20 $\times1024$ tiles from Terrain Diffusion, all from the same seed used for Fig.  1. The model produces sharp ridges, coherent river basins, smooth transitions, and varied landscapes. No visible tiling artifacts confirm the effectiveness of InfiniteDiffusion.
To demonstrate practical use, we integrate Terrain Diffusion into the Minecraft engine by replacing the native world generator. Elevation and biome queries are routed through our model, and climatic outputs are mapped to Minecraft biomes using a lightweight rule set. The system streams terrain in real time and handles arbitrary traversal; only features that rely on long distance biome searches, such as /locate biome and explorer maps, remain unsupported. Runtime is dominated by Minecraft’s own generation logic rather than our model, and gameplay remains smooth even under rapid movement. Figure 6 shows representative in game terrain. For these interactive visualizations, we apply bilinear interpolation to upsample the heightmaps $4\times$.
7. Discussion
Coherence, hydrology, and diversity.
Terrain Diffusion generates only elevation plus a few coarse climate variables, yet the resulting terrain is highly coherent. Landscapes aligns with climate, valleys reflect plausible erosion patterns, with hydrological consistency that far exceeds procedural noise, although isolated basins still occur. Qualitatively, the model synthesizes a wide range of landforms, from mountain ranges and volcanoes to canyons, fjords, and coastal plains, all without class conditioning. This diversity, combined with seamless infinite generation and real-time inference, shows that diffusion models can capture structured planetary complexity while retaining many or all the important properties of procedural noise.
Limitations.
The main limitation of our method is that each model relies on conditioning from the level above it, and global coherence deteriorates if this sequence is broken. Some top level must therefore be specified externally. In this work, simple Perlin noise is sufficient because continental scale structure is extremely coarse and contains only broad gradients, so Perlin provides a reasonable starting point while remaining seed consistent and controllable. In domains where a procedural prior is not available, the appropriate replacement is less obvious. A learned generator such as InfinityGAN, or any padding-free GAN, may provide a viable solution, since seed consistency is preserved and the simplicity of the distribution at very coarse scales makes the usual drawbacks of GANs much less significant. We experimented with an end-to-end hierarchy using a small padding-free GAN as the base model, but its outputs were largely comparable to Perlin noise and offered reduced controllability. This direction remains open for future work.
Why InfiniteDiffusion works at low $T$.
While InfiniteDiffusion is practically limited to small $T$, we find that the method retains striking coherence even when generation is aggressively truncated to $T=1$ or $2$. We hypothesize that because global structure is largely determined from hierarchical conditioning, overlapping windows only need to enforce local consistency that is already largely aligned through shared noise. Additionally, because the diffusion models are trained on random crops, they learn an approximately translation invariant representation that bases predictions on local patterns, which are shared across overlapping windows, rather than absolute position, further reducing seams. Together, these factors make the method robust even under minimal iterative refinement.
8. Future Work
Adding features to the hierarchy is a natural next step. The coarse model, the base model, or both could incorporate additional variables such as soil properties, other climatic variables, or satellite imagery, enhancing control and enabling additional downstream applications. Resolution can also be extended beyond the current hierarchy by adding further refinement stages. While increased resolution does require more work for the same real-world area, traversal speed and viewing distance usually scale with resolution. In this case, the lower throughput required at low resolutions typically balances the extra work at high resolutions, making high-resolution generation highly efficient. Finally, the InfiniteDiffusion formulation itself is not specific to terrain. Any domain that can be decomposed into overlapping tiles can adopt the same sampling strategy, including textures, maps, and large environments in general.
9. Conclusion
We have presented Terrain Diffusion, a diffusion-based framework for coherent, real-time terrain generation across planetary scales. By reformulating MultiDiffusion for unbounded domains and introducing the Infinite Tensor framework, we enable seed-consistent, random-access synthesis of infinite terrain with constant memory and compute. A hierarchical stack of diffusion and consistency models unifies planetary organization with local realism, producing continuous landscapes far beyond the reach of procedural noise. Together, these components position diffusion models as a practical foundation for procedural worldbuilding.
Appendix A Formal Properties of InfiniteDiffusion
A.1. Preliminaries and Notation
We first make precise the InfiniteDiffusion framework used in the main text.
Infinite image space.
Let the spatial domain be the integer lattice $\mathbb{Z}^{d}$, and let
$\mathcal{J}\coloneqq\mathbb{R}^{\mathbb{Z}^{d}\times C}$
denote the space of infinite images with $C$ channels. We write $J\in\mathcal{J}$ as a function $J:\mathbb{Z}^{d}\to\mathbb{R}^{C}$ over pixel coordinates. For the following sections, we write $\Omega$ as shorthand for the index-set $H_{1}]\times\cdots\times[H_{d}]$.
Regions and windows.
A (rectangular) region is any subset of the form
$R=[a_{1},b_{1})\times\cdots\times[a_{d},b_{d})\subset\mathbb{Z}^{d}$
with integers $a_{k}<b_{k}$. The set of window indices is a countable set $S$ (e.g. $S=\mathbb{Z}^{d}$). For each $i\in S$ we are given a window region $R_{i}\subset\mathbb{Z}^{d}$, a weight map $W_{i}\in\mathbb{R}^{\Omega}$, and a conditioning vector $y_{i}$. We assume the window layout is such that for every finite region $R$, only finitely many windows intersect $R$:
Assumption 1 (Finite window overlap). For every finite region $R$,
$\kappa(R)\coloneqq\{i\in S\;:\;R_{i}\cap R\neq\emptyset\}$
is finite.
Embedding operator.
For any tensor $x\in\mathbb{R}^{\Omega\times C}$, the operator $U_{i}(x)\in\mathcal{J}$ denotes the infinite image obtained by placing $x$ on $R_{i}$ and zero elsewhere.
Pretrained diffusion model.
Let $\Phi$ denote a fixed pretrained denoising network acting on window-sized images. We formalize it as a deterministic function
$\Phi:\mathbb{R}^{\Omega\times C}\times\mathcal{Y}\to\mathbb{R}^{\Omega\times C},$
where $\mathcal{Y}$ is the space of conditioning vectors.
InfiniteDiffusion update.
For a given noisy image $J_{t}\in\mathcal{J}$, the InfiniteDiffusion update at step $t\to t-1$ is defined, for any finite region $R$, by
(3) $\Psi(J_{t}\mid z)[R]=\left(\frac{\sum_{i\in\kappa(R)}U_{i}\bigl(W_{i}\otimes\Phi(J_{t}[R_{i}]\mid y_{i})\bigr)}{\sum_{j\in\kappa(R)}U_{j}(W_{j})}\right)[R],$
where $\otimes$ denotes the Hadamard (elementwise) product and the division is also elementwise. For this update, we adopt the convention that any division by zero is defined to be zero. By Assumption 1, both sums are finite on any finite $R$, so (3) is well-defined. $z$ is a (possibly infinite) vector from which the $y_{i}$ are computed.
Seeds and initial noise.
Let $\mathcal{X}$ be a set of seeds. A seed $s\in\mathcal{X}$ deterministically selects an initial noise field $J_{T}^{(s)}\in\mathcal{J}$, and a conditioning vector $z^{(s)}\in\mathcal{Z}$. Formally, there are deterministic functions
$G:\mathcal{X}\times\mathbb{Z}^{d}\to\mathbb{R}^{C},\qquad H:\mathcal{X}\to\mathcal{Z},\qquad\Lambda:\mathcal{Z}\times S\to\mathcal{Y}$
such that
$J_{T}^{(s)}(p)=G(s,p),\qquad z^{(s)}=H(s),\qquad y_{i}^{(s)}=\Lambda(z^{(s)},i).$
Definition of $J_{t}^{(s)}$.
For a fixed seed $s$, we define the entire trajectory $(J_{t}^{(s)})_{t=0}^{T}$ recursively by:
  • $J_{T}^{(s)}$ is given by $G$.
  • For $t=T,T-1,\dots,1$, define $J_{t-1}^{(s)}$ via (3) with $J_{t}=J_{t}^{(s)}$ and $y_{i}=y_{i}^{(s)}$.
Because $T$ is finite and each update uses only finite sums on any finite region, the tensors $J_{t}^{(s)}$ are well-defined for all $t\in\{0,\dots,T\}$.
Lazy evaluation algorithm.
Algorithm 1 efficiently computes (3). At level $t$ it maintains images $A_{t-1},B_{t-1}\in\mathcal{J}$ and a processed set $P_{t-1}\subseteq S$. To answer a query $J_{t-1}[R]$, it performs:
For each window $i\in\kappa(R)\setminus P_{t-1}$:
$A_{t-1}[R_{i}]$ $\leftarrow A_{t-1}[R_{i}]+W_{i}\otimes\Phi(J_{t}[R_{i}],y_{i}),$
$B_{t-1}[R_{i}]$ $\leftarrow B_{t-1}[R_{i}]+W_{i}.$
Add each such $i$ to $P_{t-1}$. The result is
$J_{t-1}[R]=A_{t-1}[R]/B_{t-1}[R].$
Where the division is performed elementwise. New windows are evaluated recursively by querying $J_{t}[\cdot]$ in the same way at level $t$, until reaching $J_{T}^{(s)}$, which is given by the seed. At intialization, $A_{t-1}=\textbf{0}$, $B_{t-1}=\textbf{0}$, and $P_{t-1}=\emptyset$.
We now formalize and prove the three properties stated in the main text.
A.2. Seed Consistency
Informally, seed consistency says that once the seed is fixed, every finite region $J_{t}[R]$ is a deterministic function of the seed and the region alone, and is independent of the order in which regions are queried.
Definition A.1 (Seed-consistent generative process).
A family of random infinite images $\{J_{t}\}_{t=0}^{T}$ on $\mathbb{Z}^{d}$ is seed-consistent if there exists a set of seeds $\mathcal{X}$ such that for all $s\in\mathcal{X}$, $t\in\{0,\dots,T\}$ and finite $R$,
$J_{t}[R]$
is a function of $s$ and $R$. Consequently, repeated queries for the same $t,s,R$ always return the same value, irrespective of the order in which regions are requested.
We show that InfiniteDiffusion, as defined in  3, is seed-consistent.
Lemma A.2 (InfiniteDiffusion is seed-consistent).
Fix a seed $s\in\mathcal{X}$. Then for each $t\in\{0,\dots,T\}$ and finite region $R$, the tensor $J_{t}^{(s)}[R]$ defined by the recursive update (3) is uniquely determined by $s$ and $R$.
Proof.
We proceed by backward induction on $t$.
Base case ($t=T$). By construction, $J_{T}^{(s)}(p)=G(s,p)$ for all $p$, so for any finite region $R$, the restriction $J_{T}^{(s)}[R]$ is uniquely determined by $s$ and $R$.
Inductive step. Assume that for some $t\in\{1,\dots,T\}$, $J_{t}^{(s)}[R]$ is uniquely determined by $s$ and $R$ for all finite regions $R$. Consider $J_{t-1}^{(s)}[R]$ for a finite region $R$.
By definition (3),
$J_{t-1}^{(s)}[R]=\left(\frac{\sum_{i\in\kappa(R)}U_{i}\bigl(W_{i}\otimes\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)})\bigr)}{\sum_{j\in\kappa(R)}U_{j}(W_{j})}\right)[R]$
For each $i\in\kappa(R)$, the region $R_{i}$ is finite, so by the inductive hypothesis $J_{t}^{(s)}[R_{i}]$ is uniquely determined by $s$ and $R_{i}$. The model $\Phi$ and weight maps $W_{i}$ are deterministic. Therefore each term
$U_{i}\bigl(W_{i}\otimes\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)})\bigr)$
is uniquely determined by $s$, and hence the finite sums in the numerator and denominator are uniquely determined. Thus $J_{t-1}^{(s)}[R]$ is uniquely determined by $s$ and $R$.
By induction, the claim holds for all $t$. ∎
Lemma A.2 shows that InfiniteDiffusion has a well-defined deterministic output for a fixed seed. We now show that the lazy query algorithm is consistent with this definition, and is therefore also seed-consistent.
Lemma A.3 (Algorithm Consistency).
Fix $s\in\mathcal{X}$ and a timestep $t\in\{1,\dots,T\}$. Then immediately before any query $J_{t-1}[R]$ with Algorithm 1, the pair $(A_{t-1},B_{t-1})$ satisfy
$A_{t-1}$ $=\sum_{i\in P_{t-1}}U_{i}\bigl(W_{i}\otimes\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)})\bigr),$
$B_{t-1}$ $=\sum_{i\in P_{t-1}}U_{i}(W_{i}).$
Furthermore, after performing one iteration of Algorithm 1, the updated state satisfies the same form with $P_{t-1}$ replaced by $P_{t-1}\cup\kappa(R)$.
Proof.
We prove this by induction over an arbitrary sequence of queries. At initialization, $A_{t-1}=0$, $B_{t-1}=0$, and $P_{t-1}=\emptyset$, so the claim is trivially true.
Now assume that before any query $J_{t-1}[R]$ we have
$A_{t-1}=\sum_{i\in P_{t-1}}U_{i}(V_{i}),\quad B_{t-1}=\sum_{i\in P_{t-1}}U_{i}(W_{i}),$
where we write $V_{i}$ as shorthand for $W_{i}\otimes\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)})$.
During the query, for each $i\in\kappa(R)\setminus P_{t-1}$ the algorithm performs
$A_{t-1}[R_{i}]\leftarrow A_{t-1}[R_{i}]+V_{i},\qquad B_{t-1}[R_{i}]\leftarrow B_{t-1}[R_{i}]+W_{i}$
and does not modify any pixels outside $R_{i}$. Equivalently, this is
$A_{t-1}\leftarrow A_{t-1}+U_{i}(V_{i}),\qquad B_{t-1}\leftarrow B_{t-1}+U_{i}(W_{i}).$
After processing all such windows we obtain
$A_{t-1}=\sum_{i\in P_{t-1}}U_{i}(V_{i})+\sum_{i\in\kappa(R)\setminus P_{t-1}}U_{i}(V_{i})=\sum_{i\in P_{t-1}\cup\kappa(R)}U_{i}(V_{i}),$
and similarly
$B_{t-1}=\sum_{i\in P_{t-1}\cup\kappa(R)}U_{i}(W_{i}).$
Finally, the algorithm updates $P_{t-1}\leftarrow P_{t-1}\cup\kappa(R)$, so the updated state satisfies the same form with $P_{t-1}$ replaced by $P_{t-1}\cup\kappa(R)$.
By induction, the claim holds for all queries $J_{t-1}[R]$. ∎
We now show that Algorithm  1 is consistent with the formal definition in  3.
Lemma A.4 (Correctness of a single query).
Fix $s$, $t$, and a finite region $R$. After any query $J_{t-1}[R]$ following Algorithm  1, we have
$J_{t-1}[R]=A_{t-1}[R]/B_{t-1}[R]=J_{t-1}^{(s)}[R].$
Proof.
By Lemma A.3, after the query finishes we have
$A_{t-1}=\sum_{i\in P^{\prime}_{t-1}}U_{i}(V_{i}),\qquad B_{t-1}=\sum_{i\in P^{\prime}_{t-1}}U_{i}(W_{i}),$
for some processed set $P^{\prime}_{t-1}\supseteq\kappa(R)$.
Now restrict to the region $R$. Any window $i\notin\kappa(R)$ has $R_{i}\cap R=\emptyset$, so
$U_{i}(V_{i})[R]=0,\qquad U_{i}(W_{i})[R]=0.$
Therefore,
$A_{t-1}[R]=\left(\sum_{i\in P^{\prime}_{t-1}}U_{i}(V_{i})\right)[R]=\left(\sum_{i\in\kappa(R)}U_{i}(V_{i})\right)[R],$
and similarly
$B_{t-1}[R]=\left(\sum_{i\in\kappa(R)}U_{i}(W_{i})\right)[R].$
Comparing with the definition (3), we see that
$\frac{A_{t-1}[R]}{B_{t-1}[R]}=\Psi(J_{t}^{(s)}\mid z^{(s)})[R]=J_{t-1}^{(s)}[R],$
which is the claim. ∎
We can now state the main seed consistency result.
Theorem A.5 (Seed consistency of the algorithm).
Under Assumption 1, the InfiniteDiffusion lazy query algorithm is seed-consistent.
Proof.
By Lemma A.2, for each $s$, $t$, and finite $R$, $J_{t}^{(s)}[R]$ is uniquely determined by $(s,R)$, so the process defined by (3) is seed-consistent. By Lemma  A.4, the algorithm is equivalent, and therefore also seed-consistent. ∎
Theorem A.5 formalizes the informal argument in the main text: once a seed is fixed, the entire trajectory $(J_{t}^{(s)})_{t=0}^{T}$ is fully determined, and the lazy querying and caching scheme merely memoizes a deterministic computation without affecting its outcome. In particular, querying regions in different orders or repeating a query for the same region cannot change the result.
A.3. Constant-Time Random Access
We now formalize the claim that accessing the value on any window-sized region has constant computational cost, independent of the absolute location in the infinite domain. We measure cost in terms of the number of evaluations of $\Phi$, which dominate runtime in practice. Let $C_{t}(R)$ denote the worst-case number of $\Phi$-calls required by the lazy algorithm to answer a query for $J_{t}[R]$, starting from an empty cache at all timesteps.
Assumption 2 (Uniform overlap bound). There exists a finite constant $M$ such that $|\kappa(R_{i})|\leq M$ for every window region $R_{i}$. In words, each window region overlaps the regions of at most $M$ windows.
We treat the total number of diffusion steps $T$ and the window shape as fixed hyperparameters of the model.
Lemma A.6 (Recursive cost bound).
Under Assumption 2, for any timestep $t$ and any window index $i\in S$, the cost $C_{t}(R_{i})$ satisfies
$C_{t}(R_{i})\;\leq\;M(1+\sup_{j\in S}C_{t+1}(R_{j}))\quad\text{for }t<T,$
with base case $C_{T}(R_{i})=0$ for all $i$.
Proof.
Consider a query for $J_{t}[R_{i}]$ at some $t<T$. By the update rule, computing this query requires evaluating $\Phi(J_{t+1}[R_{k}],y_{k})$ for every $k\in\kappa(R_{i})$.
Each such evaluation requires one call to $\Phi$ at level $t$, and in order to provide the input $J_{t+1}[R_{k}]$, the algorithm may in turn need to perform some number of step-$(t+1)$ queries. That is, for each $k\in\kappa(R_{i})$ we incur cost $1+C_{t+1}(R_{k})$.
The cardinality of $\kappa(R_{i})$ is at most $M$ by Assumption 2. Thus
$C_{t}(R_{i})\leq\sum_{k\in\kappa(R_{i})}(1+C_{t+1}(R_{k}))\leq M(1+\sup_{j\in S}C_{t+1}(R_{j}))$
At the top level $t=T$, no further calls to $\Phi$ are required because $J_{T}$ is given directly by the noise generator, hence $C_{T}(R_{i})=0$. ∎
Theorem A.7 (Uniform bound on cost for window regions).
Under Assumption 2, there exists a constant $K$ depending only on $T$ and $M$ such that for all $t$ and all window indices $i$,
$C_{t}(R_{i})\leq K.$
Proof.
Let
$c_{t}\coloneqq\sup_{i\in S}C_{t}(R_{i}).$
By Lemma A.6,
$c_{t}\leq M(1+c_{t+1}),\quad t<T,$
with $c_{T}=0$.
Unwinding this recurrence yields
$c_{T-1}\leq M(1+0)=M,\quad c_{T-2}\leq M(1+c_{T-1})\leq M+Mc_{T-1},$
and in general
$c_{t}\leq M+M^{2}+\dots+M^{T-t}.$
For fixed $T$ and $M$, the right-hand side is a finite constant independent of $i$ and the absolute location of $R_{i}$ in the plane. Taking
$K\coloneqq M+M^{2}+\dots+M^{T}$
gives the claimed uniform bound. ∎
For any window index $i$, Theorem A.7 gives a uniform bound $C_{0}(R_{i})\leq K$ on the number of calls to $\Phi$ needed to answer a query for $J_{0}[R_{i}]$, starting from an empty cache. This bound depends only on $T$ and $M$, which are constants, and not on $i$ or any regions previously processed. In standard algorithmic notation, this means that the time complexity of a query $J_{0}[R]$ is $O(1)$, when $R$ is a window region.
When caches are reused across multiple queries, subsequent queries typically cost far less than $K$, but this is not required for the asymptotic guarantee.
Combined with seed consistency, Theorem  A.7 formally justifies the claim that users can jump to arbitrary locations in the infinite world and query any window region efficiently, without needing to generate intermediate tiles and without affecting the content of the regions. When we assume that any region of fixed size has a bounded number of window overlaps, the same argument applies to arbitrary regions of bounded size, not just individual windows.
A.4. Parallelization
Finally, we formalize the statement that InfiniteDiffusion admits parallel evaluation of window updates at any fixed timestep.
Recall that, at fixed $t$ and seed $s$, the numerator and denominator images for the update $J_{t}^{(s)}\mapsto J_{t-1}^{(s)}$ are
$A_{t-1}^{(s)}$ $=\sum_{i\in S}U_{i}\bigl(W_{i}\otimes\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)})\bigr),$
$B_{t-1}^{(s)}$ $=\sum_{i\in S}U_{i}(W_{i}),$
The updated image is then
$J_{t-1}^{(s)}=A_{t-1}^{(s)}/B_{t-1}^{(s)}.$
We treat calls to $\Phi$ as the only expensive operation, and all other operations (additions, multiplications, divisions and updates to $A_{t-1},B_{t-1}$) as free. A computation is parallelizable if there exists an algorithm in which all calls to $\Phi$ can be partitioned into finitely many rounds so that within each round the calls are independent and can be executed simultaneously.
We first note that, at a fixed timestep, once the inputs $J_{t}^{(s)}[R_{i}]$ are known, all required model evaluations can be done in parallel.
Lemma A.8 (Parallel window updates at a fixed timestep).
Fix a seed $s$, a timestep $t\in\{1,\dots,T\}$, and a finite set of window indices $I\subseteq S$. Suppose that for every $i\in I$ the tensor $J_{t}^{(s)}[R_{i}]$ is available for free. Then the evaluations
$\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)}),\qquad i\in I,$
can be performed in a single parallel round, and all corresponding contributions to $A_{t-1}^{(s)}$ and $B_{t-1}^{(s)}$ on $\bigcup_{i\in I}R_{i}$ can be formed without any further calls to $\Phi$.
Proof.
For each $i\in I$, the input to $\Phi$ depends only on $J_{t}^{(s)}[R_{i}]$ and $y_{i}^{(s)}$, which are both already available. Thus the evaluations $\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)})$ are mutually independent and can be carried out simultaneously.
Once these outputs are known, the updates
$A_{t-1}^{(s)}$ $\leftarrow A_{t-1}^{(s)}+U_{i}\bigl(W_{i}\otimes\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)})\bigr)$
$B_{t-1}^{(s)}$ $\leftarrow B_{t-1}^{(s)}+U_{i}(W_{i})$
and the division $J_{t-1}^{(s)}=A_{t-1}^{(s)}/B_{t-1}^{(s)}$ involve only element-wise arithmetic and therefore require no additional model evaluations. Hence all work associated with the windows in $I$ can be completed using a single parallel round of calls to $\Phi$. ∎
We now show that answering any finite collection of region queries at any timestep admits a parallel schedule of model evaluations.
Theorem A.9 (Parallelization of finite query sets).
Fix a seed $s$, a timestep $t\in\{0,\dots,T\}$, and a finite collection of regions
$\mathcal{R}=\{R^{(1)},\dots,R^{(m)}\}.$
Consider the computation that evaluates $J_{t}^{(s)}[R^{(k)}]$ for all $k=1,\dots,m$ using the recursive update (3), starting from $J_{T}^{(s)}$ and without caching. Then all calls to $\Phi$ required by this computation can be arranged into at most TtT-t parallel rounds.
Proof.
By Lemma A.2, once $s$ is fixed the values $J_{u}^{(s)}[R]$ are uniquely determined for all $u$ and all finite regions $R$. In particular, the set of model evaluations that appear in the recursive computation of $\{J_{t}^{(s)}[R^{(k)}]\}_{k=1}^{m}$ is fixed; only their order of execution is not.
We prove the claim by backward induction on $t$.
Base case ($t=T$). By definition, $J_{T}^{(s)}$ is given directly by the noise generator $G$ and does not require any calls to $\Phi$. Thus any finite set of queries at $t=T$ is trivially parallelizable with zero rounds.
Inductive step. Fix $t<T$ and assume the statement holds for $t+1$. Consider a finite collection $\mathcal{R}=\{R^{(1)},\dots,R^{(m)}\}$ of regions at time $t$.
Let
$R^{\ast}\coloneqq\bigcup_{k=1}^{m}R^{(k)}$
be the union of all queried regions at level $t$. By Assumption 1, only finitely many windows intersect $R^{\ast}$, so the index set
$I_{t}\coloneqq\kappa(R^{\ast})=\{i\in S:R_{i}\cap R^{\ast}\neq\emptyset\}$
is finite. By (3), computing $J_{t}^{(s)}[R^{(k)}]$ for all $k$ requires knowing the regions $J_{t+1}^{(s)}[R_{i}]$ for all $i\in I_{t}$.
Each region $R_{i}$ is finite, and from the recursion defining $J_{t}^{(s)}$ we see that $J_{t}^{(s)}[R_{i}]$ itself is obtained from regions of the form $J_{t+1}^{(s)}[R_{j}]$ for windows $j$ that intersect $R_{i}$. Let $\mathcal{R}_{t+1}$ denote the (finite) collection of all such window regions $R_{j}$ at timestep $t+1$ that are needed in this way. By the induction hypothesis applied at level $t+1$ to the finite set $\mathcal{R}_{t+1}$, all model evaluations needed to compute $\{J_{t+1}^{(s)}[R_{j}]:R_{j}\in\mathcal{R}_{t+1}\}$ can be scheduled in at most T(t+1)T-(t+1) parallel rounds.
Once all these regions are available, the corresponding values $J_{t}^{(s)}[R_{i}]$ for $i\in I_{t}$ are determined and can be formed without further calls to $\Phi$. At this point, Lemma A.8 implies that all remaining evaluations
$\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)}),\qquad i\in I_{t},$
required to construct $A_{t-1}^{(s)}$ and $B_{t-1}^{(s)}$ on $R^{\ast}$ can be carried out in a single additional parallel round. No more model evaluations are needed to extract $J_{t}^{(s)}[R^{(k)}]$ from these images.
Thus, the entire computation for the queries $\mathcal{R}$ at timestep $t$ can be performed using at most
1+(T(t+1))=Tt1+(T-(t+1))=T-t
parallel rounds of calls to $\Phi$. This completes the induction. ∎
Theorem A.9 shows that for any fixed seed $s$, timestep $t$ and finite collection of regions $\mathcal{R}$, the recursive computation of $\{J_{t}^{(s)}[R]:R\in\mathcal{R}\}$ admits a bounded-depth parallel schedule of diffusion model evaluations. Combined with seed consistency, this formalizes the claim in the main text that all diffusion model evaluations required to serve any finite batch of queries $J_{t}^{(s)}[R]$ can be performed in parallel, up to the intrinsic sequential dependence across diffusion steps.